# FIG: Forward-Inverse Generation for Low-Resource Domain-specific Event Detection

### v1

@@ Abstract
Event Detection (ED) is the task of identifying typed event mentions of interest from natural language text, which benefits domain-specific reasoning in biomedical, legal, and epidemiological domains. However, procuring supervised data for thousands of events for various domains is a laborious and expensive task. To this end, existing works have explored synthetic data generation via forward (generating labels for unlabeled sentences) and inverse (generating sentences from generated labels) generations. However, forward generation often produces noisy labels, while inverse generation struggles with domain drift and incomplete event annotations. To address these challenges, we introduce FIG, a hybrid approach that leverages inverse generation for high-quality data synthesis while anchoring it to domain-specific cues extracted via forward generation on unlabeled target data. FIG further enhances its synthetic data by adding missing annotations through forward generation-based refinement. Experimentation on three ED datasets from diverse domains reveals that FIG outperforms the best baseline, achieving average gains of 3.3% F1 and 5.4% F1 in the zero-shot and few-shot settings, respectively. Analyzing the generated trigger hit rate and human evaluation substantiates FIG’s superior domain alignment and data quality compared to existing baselines. We will release our code at https://github.com/PlusLabNLP/FIG.
FIG: Forward-Inverse Generation for Low-Resource
Domain-specific Event Detection
@@ 1 Introduction
Figure 1: Comparing different data generation paradigms. Owing to poor LLM ED reasoning skills, forward generation (top) struggles with noisy labels. Inverse generation (middle) suffers from domain drift owing to lack of domain-specific cues. We introduce FIG (bottom), a hybrid forward and inverse approach, to overcome both issues to generate better quality data.
Event Detection (ED) Sundheim (1992); Doddington et al. (2004) involves identifying and categorizing significant events from natural language text based on a pre-defined ontology. It has applications in widespread domains like biomedicine Pyysalo et al. (2012), epidemiology Parekh et al. (2024b, c), law Francesconi et al. (2010), etc. However, procuring annotations for each domain-specific ontology to train models is expensive and impractical. Although recent works introduce zero-shot LLM-reasoning Gao et al. (2023); Cai et al. (2024), their performance falls short of supervised approaches Huang et al. (2024).
To mitigate the need for extensive annotations to train supervised models, prior work has explored LLM-powered synthetic data generation to enhance downstream model training. A widely used approach is a forward generation or weak supervision He et al. (2021); Chia et al. (2022), where labels are assigned to existing unlabeled data (i.e., $X\rightarrow Y^{\prime}$). This generation is limited by LLM ED reasoning, often leading to noisy labels. To address this, inverse generation has been proposed Meng et al. (2022); Wang et al. (2023c), where sentences are generated based on sampled or generated labels (i.e $Y\rightarrow X^{\prime}$). However, inverse generation introduces a domain drift between synthetic and target data owing to the lack of any domain-specific cues, as illustrated in Figure 1.
In our work, we propose Forward-Inverse Generation (FIG), a hybrid approach that integrates forward and inverse generation to enhance synthetic data quality. Instead of generating event triggers from scratch — an inherently high-variance process — we first apply forward generation to unlabeled data, extracting domain-specific triggers. Next, we utilize inverse generation to synthesize diverse sentences conditioned on the domain-specific triggers. Finally, a second forward generation annotates any missing events in the generated data to ensure high data quality. This approach produces cleaner synthetic data compared to forward generation while maintaining a closer domain alignment with target data than inverse generation, as shown in Figure 1.
We evaluate FIG on ED datasets from three different domains: ACE Doddington et al. (2004) (news), SPEED Parekh et al. (2024c) (social media), and GENIA2011 Kim et al. (2011) (biomedical). Our primary method of evaluation is testing ED performance of DEGREE Hsu et al. (2022) trained on the generated data. In the zero-shot setting, FIG outperforms STAR Ma et al. (2024) (inverse generation baseline) and weak supervision (forward generation baseline) by an average of 16.8% and 3.5% Tri-C F1, respectively. Similarly, in the few-shot setting, FIG surpasses the strongest baseline by an average of 6.2% Tri-C F1. We study the generated trigger hit rate (relative to the gold trigger set) and demonstrate how FIG’s triggers have about 7.3% better hit rate compared to the best baseline, which contributes to its superior performance. Finally, we also conduct human evaluation, which supports FIG’s superior domain relevance and data annotation quality.
@@ 2 Problem Definition
We focus on the task of Event Detection (ED) Sundheim (1992); Doddington et al. (2004) for this work. The task of ED aims to extract mentions of any events of interest from natural language text. Following ACE 2005 Doddington et al. (2004), we define an event as something that happens or describes a change of state and is labeled by a specific event type. The word/phrase that most distinctly highlights the occurrence of the event is defined as the event trigger, and the trigger-event type pair is referred to as the event mention. Event Detection, in particular, requires extracting the event triggers from the sentence and classifying it into one of the pre-defined event types. We provide an illustration of this task below where arrested and campaigns are the trigger words for the event types Justice: Arrest-Jail and Conflict: Demonstrate respectively.
@@
In our work, we specifically focus on ED in diverse and specialized domains (e.g., biomedical), where procuring a training dataset $D_{T}$ of annotated data points is expensive, whereas reasonable-sized unlabeled data $D_{T}^{\prime}$ is available. We focus on two realistic low-resource data setups - zero-shot (zero labeled data) and few-shot ($k$ labeled datapoints per event type) settings. Different from domain transfer, we do not consider any labeled data for the source domain and directly optimize model performance for the target domain.
@@
Figure 2: Model Architecture Diagram. Top left - illustration of forward generation. Bottom left - illustration of inverse generation. Right - FIG and its four components. We first extract domain-specific triggers via forward generation, then generate passages using inverse generation. Forward generation refines missing events, and we sample $N$ data points per event for downstream training.
3 Methodology
In this work, we focus on Large Language Model (LLM)-powered synthetic data generation to alleviate the need for domain-specific annotated training data, as zero-shot LLM-based approaches usually struggle in ED in specialized domains Huang et al. (2024). By generating a large amount of data instances $D_{s}=\{(X,Y)\}$, we can train downstream supervised ED models. Here, we first briefly review prior works on forward generation and backward generation. Next, we introduce our proposed data generation method FIG.
3.1 Background
We focus on two major paradigms for data generation in our work: forward generation and inverse generation. We describe them briefly below and provide illustrations in Figure 2.
Forward Generation:
This is a more straightforward manner of data generation, wherein LLMs are utilized to generate labels $Y^{\prime}$ on unlabeled data $X\in D_{T}^{\prime}$ (i.e. $X\rightarrow Y^{\prime})$. This generation is analogous to weak/distant supervision (Weak Sup) Mintz et al. (2009); Wang et al. (2021); He et al. (2021) where noisy/weak labels are assigned to clean sentences and then utilized to train downstream models. Consequently, the label quality is dependent on the reasoning capability of the LLM.
Inverse Generation:
Inverse generation is a relatively new data generation paradigm Kumar et al. (2020); Schick and Schütze (2021). This paradigm first generates/samples potential labels $Y$ based on the task definition and then generates plausible $X^{\prime}$ conditioned on the label $Y$ (i.e. $Y\rightarrow X^{\prime}$). Recent works like SynthIE Josifoski et al. (2023) and STAR Ma et al. (2024) have explored the usage of LLM-guided inverse generation for information extraction and event extraction tasks. Inverse generation provides control over data distribution via prompting, but the data quality is dependent on the LLM’s sentence generation capability.
3.2 FIG
LLM reasoning has been poor for ED Li et al. (2023a); Huang et al. (2024), which makes forward generation vulnerable to noisy label quality. Since LLMs are pre-trained to generate natural sentences, their generation capabilities are stronger, which supports inverse generation. However, owing to the lack of any domain-specific information, inverse generation can synthesize highly diverse sentences, leading to a domain drift. Merely specifying the domain in the prompt is not sufficient (shown in § D.1). Furthermore, inverse generation sentences can mention additional events which remain unannotated introducing noise in the data.
To this end, we propose Forward-Inverse Generation (FIG), a hybrid approach that infuses forward generation reasoning into the base pipeline of inverse generation to ensure closer alignment to the target domain while improving the quality of the annotated labels. To procure domain-specific cues, we assume access to an unlabeled target domain data $D_{T}^{\prime}$ (similar to forward generation). We provide our architectural diagram in Figure 2 and explain each component of our pipeline below.
Target Structure Extraction:
Similar to inverse generation, we first create potential triggers (i.e. labels) for each event type. However, generation from scratch can lead to very divergent set of triggers. For example, an “Attack" event can refer to a war in the news, cyberattacks in the cybersecurity domain, diseases in epidemiology, or criticism in economics. Naturally, each scenario would assume different triggers, which inverse generation struggles to generate even when provided with clear event definitions (actual examples in § 6.5).
@@ 3.2 FIG
Instead, we generate potential triggers per event type by utilizing an LLM to extract triggers using a forward generation over the unlabeled data $D_{T}^{\prime}$. To ensure highly precise extraction of triggers, we develop a two-stage prompt setup. The first stage is tasked to identify and filter possible event types mentioned in the text based on the task instructions and event definitions. The second stage aims to find the most appropriate trigger word from the unlabeled sentence for each filtered event type. We illustrate this setup in § A.
After extraction, we aggregate and sort the triggers for each event type at the corpus level and filter out the top $t$ triggers for each event type as the set of clean triggers. This filtering helps remove many noisy triggers introduced in the forward generation. Finally, we create target structures $Y$ by sampling 1-2 event types per instance and sampling triggers from the clean trigger set for each sampled event type. We ensure uniform sampling of event types and triggers to avoid unbalanced label distribution. Overall, such forward generation-based trigger extraction ensures the distillation of domain-specific knowledge in the target structures.
Passage Generation:
This component is the core of inverse generation and is tasked with generating passages corresponding to the constraints of the sampled target structure. Specifically, the LLM is prompted with the task instructions, the event definitions, the sampled target structure $Y$, and asked to generate a passage $X^{\prime}$ which mentions the event types using the triggers in $Y$. We illustrate this prompt in Figure 8. The key to generating better target data aligned passages lies in the presence of the target data aligned target structure $Y$ (shown via examples in § 6.5).
Data Refinement and Sampling:
While passage generation ensures the target structure events are mentioned in the passage, the passage could still have other unknown events which can lead to under-annotated $(X^{\prime},Y)$ data instances. We provide an illustration in Figure 3 with the target trigger positive, but the generated sentence also has a missing mention of symptom event triggered by got. To account for such missing annotations, we utilize forward generation to identify all possible event mentions in the generated passage. We remove duplicates and add the new event mentions to create a more complete target structure $Y_{f}$.
To further improve data quality, we apply an automated rule to remove passages that do not mention the target trigger. Additionally, we standardize trigger annotations by correcting variations in trigger word forms. Finally, we apply a greedy sampling algorithm to sample $N$ instances $(X^{\prime},Y_{f})$ per event type to create our final synthetic dataset $D_{s}$.
Downstream Model Training:
The final component utilizes the generated synthetic data $D_{s}$ to train downstream ED models in a supervised manner. The trained ED models are then used for inferring on the test set and for eventual evaluation.
4 Experimental Setup
@@
FIG(ours) train 57.4 50.2 44.6 31.5 35.2 28.9 45.7 36.9
FIG (ours) external 57.7 52.6 47.8 32.9 33.6 24.6 46.4 36.7
@@
FIG (ours) train 58.1 53.8 49.9 38.7 38.0 29.7 48.7 40.7
FIG (ours) external 59.7 55.6 50.1 39.2 39.2 31.5 49.7 42.1
@@
FIG (ours) train 54.8 48.3 50.3 36.8 39.3 31.1 48.1 38.7
FIG (ours) external 54.0 48.5 50.1 36.1 38.7 29.4 47.6 38.0
- Gold Data - 64.6 61.6 64.0 53.5 51.3 44.0 60.0 53.0
Table 1: Zero-shot results comparing FIG with other baselines across three datasets and three base LLMs. Except for Inference, all other evaluations are performances of downstream DEGREE Hsu et al. (2022) model trained on data generated by each technique. Eve-I: Event Identification F1, Tri-C: Trigger Classification F1.

### v2

# SNaRe [Uncaptioned image]: Domain-aware Data Generation for Low-Resource Event Detection

@@ Abstract
Event Detection (ED) – the task of identifying event mentions from natural language text – is critical for enabling reasoning in highly specialized domains such as biomedicine, law, and epidemiology. Data generation has proven to be effective in broadening its utility to wider applications without requiring expensive expert annotations. However, when existing generation approaches are applied to specialized domains, they struggle with label noise, where annotations are incorrect, and domain drift, characterized by a distributional mismatch between generated sentences and the target domain. To address these issues, we introduce SNaRe, a domain-aware synthetic data generation framework composed of three components: Scout, Narrator, and Refiner. Scout extracts triggers from unlabeled target domain data and curates a high-quality domain-specific trigger list using corpus-level statistics to mitigate domain drift. Narrator, conditioned on these triggers, generates high-quality domain-aligned sentences, and Refiner identifies additional event mentions, ensuring high annotation quality. Experimentation on three diverse domain ED datasets reveals how SNaRe outperforms the best baseline, achieving average F1 gains of 3-7% in the zero-shot/few-shot settings and 4-20% F1 improvement for multilingual generation. Analyzing the generated trigger hit rate and human evaluation substantiates SNaRe’s stronger annotation quality and reduced domain drift. We will release our code at https://github.com/PlusLabNLP/SNaRe.
SNaRe [Uncaptioned image]: Domain-aware Data Generation for
Low-Resource Event Detection
@@ 1 Introduction
Figure 1: Highlighting the errors of existing data generation approaches. (a) Using LLMs to generate the labels from in-domain sentences leads to label noise owing to poor LLM reasoning. (b) Utilizing LLMs to generate sentences conditioned on event and domain causes domain drift, wherein the synthetic sentence is not aligned with the target domain. Finally, in (c), we illustrate how SNaRe minimizes both errors to generate higher quality synthetic data.
Event Detection (ED) Sundheim (1992); Doddington et al. (2004) involves identifying and categorizing significant events from natural language text based on a pre-defined ontology. It has widespread applications in domains such as biomedicine Pyysalo et al. (2012), epidemiology Parekh et al. (2024b, c), law Francesconi et al. (2010). Due to the high cost of expert-annotated data, synthetic data generation Xu et al. (2023) (i.e., generating sentence and event annotations) has emerged as a promising alternative, particularly for practical use-cases in specialized domains.
However, existing generation approaches often focus on general-domain settings and fail to address the distinct challenges of specialized domains Song et al. (2025). Weak supervision methods He et al. (2021); Chia et al. (2022) that use LLMs to generate labels for unlabeled sentences frequently introduce label noise (Figure 1(a)), where incorrect or incomplete labels arise due to weak LLM reasoning Huang et al. (2024) or limited domain knowledge Song et al. (2025). Downstream training on such incorrect labels can cause spurious bias propagation. Conversely, recent generation approaches Josifoski et al. (2023); Ma et al. (2024) that utilize LLMs’ self-knowledge to jointly generate labels and sentences struggle with domain drift (Figure 1(b)), often synthesizing sentences that are misaligned with the target domain. This can be attributed to the lack of utilization of target domain information for generation and can drastically hamper model training as lexical and structural cues are highly influential for ED Tong et al. (2022). Overall, label noise and domain drift reduce synthetic data quality, eventually leading to subpar supervised downstream model performance.
To this end, we propose SNaRe [Uncaptioned image], a novel domain-aware, three-stage data synthesis LLM pipeline comprising the Scout, Narrator, and Refiner modules. Scout surveys unlabeled target domain data to identify salient triggers via prompt-based trigger extraction. Using corpus-level statistics for automated aggregation and filtering, Scout curates a list of high-quality domain-specific triggers per event type. Next, Narrator samples from these domain-specific triggers and utilizes LLMs to synthesize diverse sentences for each event type. Utilizing specialized and diverse domain information for conditional generation aids Narrator in generating more domain-aligned sentences, eventually reducing domain drift in the synthesized sentences. Since Narrator sentences could mention additional events apart from the input set, we design the Refiner to utilize LLM inference to annotate such missing events in these sentences. Narrator’s conditional text generation and Refiner’s missing label annotation aid in reducing the label noise and ensuring high data quality. We provide an illustration of SNaRe’s generation in Figure 1(c).
We benchmark SNaRe on ED datasets from three domains: ACE Doddington et al. (2004) (news), SPEED Parekh et al. (2024c) (epidemiology), and GENIA2011 Kim et al. (2011) (biomedical). For evaluation, we report the ED performance of DEGREE Hsu et al. (2022) trained on the synthesized data. Across the zero-shot and few-shot settings, SNaRe performs the best, outperforming the previous state-of-the-art baselines Ding et al. (2023); Ma et al. (2024) by an average of 3-7% F1 points. Under multilingual generation for Arabic and Chinese, SNaRe outshines even more, with improvements of 4-20% F1 over the best baseline. Our analysis reveals how SNaRe’s synthesized triggers overlap 4-11% more (relative to baselines) with the gold trigger set, demonstrating the reduction in domain drift. Finally, human evaluation provides qualitative evidence for SNaRe’s superior data annotation quality and domain alignment. [^1] We will release our code and data upon acceptance.
@@ 2 Problem Definition
We focus on the task of Event Detection (ED) Sundheim (1992); Doddington et al. (2004) for this work. ED aims to extract mentions of any events of interest from natural language text. Following ACE 2005 Doddington et al. (2004), we define an event as something that happens or describes a change of state and is labeled by a specific event type. The word/phrase that most distinctly highlights the occurrence of the event is defined as the event trigger, and the trigger-event type pair is known as the event mention. Event Detection requires extracting the event triggers from the sentence and classifying them into one of the pre-defined event types. We provide an illustration of this task below, where arrested and campaigns trigger the events of Justice: Arrest-Jail and Conflict: Demonstrate, respectively.
@@
In our work, we specifically focus on ED in diverse and specialized domains (e.g., biomedical), where procuring a training dataset $D_{T}$ of annotated data points is expensive, but unlabeled data $D_{T}^{\prime}$ is available. We focus on two realistic low-resource data setups - zero-shot (zero labeled data) and few-shot ($k$ labeled datapoints per event type) settings. Unlike domain transfer, we do not consider any labeled data for the source domain, and directly optimize model performance for the target domain.
3 Related Works
@@
Figure 2: Model Architecture Diagram highlighting the various components of SNaRe. First, Scout extracts and filters domain-specific triggers, then Narrator generates passages conditioned on these triggers. Finally, Refiner adds any missing annotations and sample $N$ data points per event for downstream training.
Data Generation for Information Extraction
LLM-powered synthetic data generation has been successful for various NLP tasks Li et al. (2023b); Wang et al. (2023c); Wu et al. (2024); Shao et al. (2025). For information extraction, works have explored knowledge retrieval Chen and Feng (2023); Amalvy et al. (2023), translation Parekh et al. (2024a); Le et al. (2024), data re-editing Lee et al. (2021); Hu et al. (2023), and label extension Zhang et al. (2024). Recent works utilize LLMs to generate labels for sentences Chia et al. (2022); Ye et al. (2022); Wang et al. (2023a); Tang et al. (2023), while some other works explore the generation of sentences from labels Josifoski et al. (2023); Ma et al. (2024). Our work introduces SNaRe focused on infusing better domain-specific information for data generation.
Low-resource Event Detection
Event Detection (ED) has been studied extensively Sundheim (1992); Grishman and Sundheim (1996), leading to diverse datasets in news Doddington et al. (2004); Song et al. (2015); Ellis et al. (2015), Wikipedia Li et al. (2021); Pouran Ben Veyseh et al. (2022), and general domains Wang et al. (2020); Parekh et al. (2023), as well as niche areas like biomedical Pyysalo et al. (2012); Kim et al. (2011, 2013), multimedia Li et al. (2020), cybersecurity Satyapanich et al. (2020), epidemiology Parekh et al. (2024b, c), and pharmacovigilance Sun et al. (2022). To address the growing need for event detection across expanding domains, prior works have explored transfer learning via Abstract Meaning Representation Huang et al. (2018), Semantic Role Labeling Zhang et al. (2021), and Question Answering Lyu et al. (2021). Reformulating ED as a conditional generation task has also aided low-resource training Hsu et al. (2022, 2023b); Huang et al. (2022). Recently, LLM-based reasoning Li et al. (2023a); Gao et al. (2023); Wang et al. (2023b) and transfer-learning Cai et al. (2024) has been explored, but their performance remains inferior to supervised models Huang et al. (2024). This motivates efforts in LLM-powered synthetic data generation for low-resource ED.
4 Methodology - SNaRe
In this work, we focus on domain-aware synthetic data generation using Large Language Models (LLMs) to alleviate the need for expert-annotated training data. By generating a large dataset $D_{s}=\{(X,Y)\}$, we can train domain-specific downstream ED models using minimal supervised data.
Existing weak-supervision approaches Mintz et al. (2009); Wang et al. (2021) utilizing automatic methods to assign labels to unlabeled sentences often generate incorrect labels (label noise). This can be attributed to the domain-specific context understanding and deep reasoning requirement of ED, leading to poor automatic label quality, even when using recent LLMs Huang et al. (2024). Another route of approaches that utilize LLMs to generate sentences conditioned on labels Schick and Schütze (2021); Josifoski et al. (2023), i.e., synthesize $X$ for a designated $Y$, often curate sentences that are distributionally divergent from the target domain (domain drift). This is mainly since these approaches focus on the general domain and fail to utilize any target domain signals in their generation. Both label noise and domain drift hurt the synthetic data quality, in turn, diminishing the downstream supervised model performance.
To mitigate these issues, we propose SNaRe [Uncaptioned image], a domain-aware high-quality data synthesizer, comprising three components: Scout, Narrator, and Refiner. Scout studies unlabeled target domain data $D_{T}^{\prime}$ to curate domain-specific triggers, in turn, reducing domain drift. Narrator generates domain-specific sentences conditioned on Scout’s curated triggers, while Refiner adds additional annotations to ensure high-quality labels. Overall, SNaRe is a training-free LLM inference pipeline and easily deployable and scalable across domains. We provide our architectural diagram in Figure 2 and explain each component of our pipeline below.
4.1 Scout
Scout is tasked with the curation of domain-specific triggers that are later utilized for sentence generation. Events can assume a wide range of triggers depending on the domain and context. For example, an "Attack" event can be triggered by war, killed in news, or breach, phish in cybersecurity, or infect, transmit in epidemiology domains.
@@ 4.1 Scout
Thus, unlike past works Ma et al. (2024) that utilize only LLMs’ internal knowledge for trigger generation, we develop Scout, which extracts high-precision domain-specific triggers using unlabeled target domain data $D_{T}^{\prime}$. Specifically, trigger extraction involves a two-stage prompt setup: (1) The first stage is tasked with identifying and filtering possible event types mentioned in the target domain sentence, and (2) The second stage aims to find the most appropriate trigger word from the unlabeled sentence for each filtered event type. We illustrate this setup in § A.
To ensure high-precision of the triggers, we develop an aggregation and filtering mechanism by incorporating corpus-level statistics. Specifically, for each event type, we aggregate the counts of the triggers at the corpus level and filter out the top $t=10$ triggers as the curated list of high-quality domain-specific event-indexed triggers. These triggers carry important target domain signals that help in generating domain-specific sentences (§ 4.2), in turn reducing domain drift of our synthetic data.
4.2 Narrator
Narrator is tasked with the synthesis of domain-specific sentences for our synthetic dataset. Existing works Josifoski et al. (2023); Ma et al. (2024) do not utilize any target domain information, which causes domain drift in their synthesized sentences. Instead, in our work, we condition the Narrator to utilize the rich and diverse domain-specific triggers from Scout to synthesize domain-specific sentences, which, in turn, reduce domain drift.
Specifically, Narrator samples 1-2 event types per synthetic data instance and corresponding domain-specific triggers from Scout’s curated trigger list – constituting the label $Y$. Next, it prompts the LLM with the task instructions and the event definitions, and asks it to generate a passage $X^{\prime}$ that mentions the sampled events using the sampled triggers ($Y$). We illustrate this prompt in Figure 8. The generated sentences are naturally more aligned with the target domain owing to the conditioning on the domain-specific triggers (qualitative examples shown in § 7.6). Further domain-specific adaptations are possible by fine-tuning the LLM on unlabeled domain-specific data (analysis in § 7.5).
@@ 4.2 Narrator
SNaRe (ours) train 57.4 50.2 44.6 31.5 35.2 28.9 45.7 36.9
SNaRe (ours) external 57.7 52.6 47.8 32.9 33.6 24.6 46.4 36.7
@@
SNaRe (ours) train 58.1 53.8 49.9 38.7 38.0 29.7 48.7 40.7
SNaRe (ours) external 59.7 55.6 50.1 39.2 39.2 31.5 49.7 42.1
@@
SNaRe (ours) train 54.8 48.3 50.3 36.8 39.3 31.1 48.1 38.7
SNaRe (ours) external 54.0 48.5 50.1 36.1 38.7 29.4 47.6 38.0
(Upper Bound) Gold Data - 64.6 61.6 64.0 53.5 51.3 44.0 60.0 53.0
Table 1: Zero-shot results comparing SNaRe with other baselines across three datasets and three base LLMs. Except for Inference, all other evaluations are performances of downstream DEGREE Hsu et al. (2022) model trained on data generated by each technique (50 datapoints per event type).
4.3 Refiner
While the Narrator ensures the sampled triggers are mentioned in the passage, it could potentially introduce new unknown events in the passage, leading to under-annotated $(X^{\prime},Y)$ data instances. We illustrate this in Figure 3 where the sampled trigger positive for infect event is mentioned, but the sentence also mentions the missing symptom event triggered by got.
To account for such missing annotations, we introduce Refiner - tasked to annotate missing events in the generated sentence. Here, we simply prompt the LLM to find all event mentions in the generated sentence (illustrated in Figure 9) and append them to the original sampled $Y$. To avoid adding noisy labels, we only add annotations for newly discovered events. To further improve data quality, we apply an automated rule to remove passages that do not mention the target trigger. Additionally, we standardize trigger annotations by correcting variations in trigger word forms. Such normalization and conservative filtering further aid in ensuring high data quality. Finally, we apply a greedy sampling algorithm to sample $N=50$ instances $(X^{\prime},Y)$ per event type to create our final synthetic dataset $D_{s}$.
Downstream Model Training:
The final component utilizes the generated synthetic data $D_{s}$ to train downstream ED models in a supervised manner. The trained ED models are then used to infer on the test set and for eventual evaluation. Since we use small BART-based language models for inference, our inference time computation is negligible compared to LLM inference methods.
5 Experimental Setup